An over-the-counter multi-vitamin has 250 mg tablets that are 15% calcium by mass. If you take 2 tablets, what mass of calcium have you ingested?
Answer #1

Answer

75mg

Explanation

Mass of one tablet = 250mg

mass % of calcium in the tablet = 15%

mass of calcium in one tablet = (15/100)×250mg = 37.5mg

mass of calcium in two tablets = 2 × 37.5mg = 75mg

Therefore,

75mg of calcium ingested